OpenStudy (anonymous):

Solve the system of equations. y = 2x² – 3 y = 3x – 1 There is no solution. (–1/2, 5), (2, –5/2) (–1/2, –5/2), (2, 5) (1/2,5/2 ), (2, 5)

OpenStudy (across):

\[2x^2-3=3x-1\]\[2x^2-3x-2=0\]\[(2x+1)(x-2)=0\]Can you finish it up?

OpenStudy (anonymous):

umm. i'll try i think it's the third one?

OpenStudy (across):

Yes. :)

OpenStudy (anonymous):

thanks